For the reaction
?Al+?Cl2 →?AlCl3 ,
what is the maximum amount of AlCl3 which could be formed from 6 mol of Al and 7.88 mol of Cl2?
Answer in units of g.
1
Expert's answer
2020-06-02T14:09:09-0400

The equation is

2Al + 3Cl2 = 2AlCl3

Chlorine is the limiting reactant, that's why the calculations of the formed product will be performed using the quantity of Cl2

n(Cl2) = 7.88 mol

n(AlCl3) = 2/3 n(Cl2) = 5.25 mol

m(AlCl3) = n(AlCl3)*M(AlCl3) = 700.04 g

The maximum amount of aluminum chloride is 700.04 g